OpenStudy (anonymous):

So in class i'm learning to add and subtract rational expressions how do I simplify once It's been factored out?! Can someone please explain, heres ex. 6 3x 1 ____ + ___ - _______ x-6 x+5 (x-6)(x+5) or [(6)/(x-6)]+[(3x)/(x+5])-[(1)/(x-6)(x-5)] Options: A. (3x+5)/(x^2-x-30) B. (3x^2-12x+29)/(x^2-x-30) C. (3x^2+6x+5)/(x^2-x-30) D. (3x^2+5)/(x-6)

OpenStudy (anonymous):

First of all you must make sure the denominator (the number under the line) are all equal

OpenStudy (anonymous):

\[\frac{ 6\color{red}{(x+5)} }{ (x-6)\color {red}{(x+5)} }+\frac{ (3x)\color{red}{(x-6)} }{ (x+5)\color{red}{(x-6)} }-\frac{ 1 }{ (x-6)(x+5) }\]

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

Now do i distribute the terms and then simplify or cancel them out?

OpenStudy (anonymous):

best to first combine them all in a single fraction

OpenStudy (anonymous):

and then distribute

OpenStudy (anonymous):

Shoot I mean 3x^2+12x+29/x^2-x-30 ? Is that correct?

OpenStudy (anonymous):

-12x, the rest is correct

OpenStudy (anonymous):

Awesome thanks!! :D I'll remember those steps!! kinda got out of order